michelle1803:

Order the set of numbers from least to greatest: 2.71 repeating 71, 2 and 3 over 4, square root 5, 5 over 2

jhonyy9:

so i think these wan be : 2,(71) \[2\frac{ 3 }{ 4 }\] \[\sqrt{5}\] \[\frac{ 5 }{ 2 }\]

jhonyy9:

so these numbers you need order from least to greatest

Mercury:

easiest method would be to convert them all to decimals ex: 2 and 3/4 is equal to 2.75, which is greater than 2.71repeating, so 2 and 3/4 is greater than 2.71repeating repeat this logic with the rest of the numbers, putting them all in order